# TeleRAG: Efficient Retrieval-Augmented Generation Inference with Lookahead Retrieval

### v1

@@ 4.4 Implementation
To enable concurrent CPU-GPU data transferring and LLM generation, we utilize PyTorch’s _copy(non_blocking=True) API and use separate CUDA streams for prefetching and LLM. In this way, the data copy operations will not block the GPU to perform computation, and thus the TeleRAG can do prefetching in parallel to the pre-retrieval LLM generation.
@@ 5 Evaluation
##\## of chunks 21 million
##\## of IVF cluster 4096
Embed model Contriever [35]
@@
Table 2: Detailed configurations of our retrieval index.
@@
We conduct extensive experiments to evaluate the effectiveness of TeleRAG. In this section, we describe the necessary details on how we set up the evaluations, present experiments results and provide in-depth analysis and discussions.
5.1 Datasets and RAG Pipelines
@@
Datastore. We build a datastore based on the wiki_dpr dataset [48], which is a popular dataset contains 2.1 billions tokens from Wikipedia. Following previous works [11, 48, 64], we chunk the passages by every 100 tokens, and use Contriever [35] to generate embedding for each chunk. The embeddings have a hidden dimension of 768.
@@
Vector index. For the baseline retrieval, we build an IVF vector index using FAISS [25] from the datastore, following setups from  [11]. See Table 2 for the detailed configurations of our vector index. As described in §4.4, we then convert the FAISS index to a customized index in PyTorch for TeleRAG.
@@
LLMs. We evaluate TeleRAG on the Llama model family [82] in three different sizes (Llama-3.2-3B, Llama-3-8B, Llama-2-13B) to represent different use cases.
RAG pipelines. We evaluate TeleRAG with six popular RAG pipelines. Figure 7 shows the overview of each pipeline. Note that even though some pipelines lack pre-retrieval generation, the post-retrieval generation serves similar functionality for the retrieval of the next iteration. Below are brief descriptions to evaluated RAG pipelines. For iterative-based pipelines, we set the maximum iterations to 3.
Figure 7: Overview of six RAG pipelines that we evaluate.
  1. 1.
    HyDE[27] prompts LLM to generate a hypothetical paragraph and perform retrieval based on the embedding of the generated paragraph.
  2. 2.
    SubQestion (SubQ) [^4] Implemented in LlamaIndex as SubQuestionQueryEngine[58] prompts LLM to generate multiple sub-questions and performs retrievals for each generated sub-question.
  3. 3.
    Iterative (Iter) [^5] Implemented in LlamaIndex as MultiStepQueryEngine[58] prompts LLM to generate narrower questions first and iteratively refine them based on previous answers. At the end of each iteration, it prompts LLM to judge if the answer is good enough.
  4. 4.
    Iter-RetGen (IRG) [78] iteratively do retrieval and LLM generation for 3 iterations.
  5. 5.
    FLARE [43] iteratively issues retrievals based on the confidence (probability score) of predicted tokens for the upcoming sentence.
  6. 6.
    Self-RAG (S-RAG) [^6] We evaluate on the short-form version which only has one iteration. [11] decides whether to retrieve documents and then answers for each query. A critique stage selects the most confident answer. We use fine-tuned models based on Llama2-7B and Llama2-13B from their official implementation [12].
Evaluation Datasets. We evaluate on three commonly used question-answering datasets, NQ [54], HotpotQA [91] and TriviaQA [47]. For each dataset, we randomly sample 512 queries and report the average unless otherwise specified.
5.2 Experiment Setups
Hardware setups. We evaluate TeleRAG on two hardware environments, Desktop and Server, which equiped with represent the settings for the desktop and cloud instance use case. For Desktop, it’s paired with a RTX 4090 GPU (24GB memory) and we consider the 3B and 8B models. For Server, it’s paired with a RTX 4090 GPU (24GB memory) and we consider the 8B and 13B models. These sizes are common model size for RAG applications [11, 10]. We do not evaluate the 13B model on Desktop as its model size (26 GB) exceeds RTX4090’s memory space. Table 3 summaries the hardware configurations.
Nprobe and top-$k$. We evaluate latency with three nprobe values, 128, 256, and 512, which represent 3.1%, 6.3%, and 12.5% of the total clusters in our evaluated index. This range falls into the commonly used settings of IVF index [106, 25]. We use a top-$k$ number of 3 for retrieval, which returns 3 most relevant passages based on the retrieval query.
RAG pipeline implementation. To collect the input and output texts for each stage and have consistent results across all pipelines, we implement the RAG pipelines with the FlashRAG framework [45]. For IRG, FLARE, and S-RAG, we use the framework’s default implementations; for the other pipelines, we re-implement them using FlashRAG’s APIs.
Benchmark methodology. We follow a benchmarking methodology from  [101] and use GPT-3.5-Turbo [69] to run through each pipeline and record the input and output text of each step. During latency evaluation, we set LLM’s output length based on the recorded text. This way, we ensure a fair latency comparison across different LLM models.
Baseline systems. To evaluate the latency of each pipeline, we construct a clean execution flow in Python that only contains LLM generation, datastore retrieval, and other necessary logical operations to fulfill each pipeline. For LLM generation, we use SGLang [101], which is a state-of-the-art LLM inference engine.
Prefetching budget setups. Based on the methodology we described in §4.3, we profile each RAG pipeline with 64 random samples from NQ [54] and define the prefetch budget of each pipeline.
Max prefetching memory. We set a maximum GPU memory limit for prefetching in each configuration. For Server with H100, we allocate up to 16GB, sufficient for a high cluster hit rate at the largest tested nprobe value (512). For Desktop with RTX4090, we allocate up to 10GB and 3.75GB for the 3B and 8B models, respectively. These settings demonstrate that TeleRAG can efficiently operate using only a small fraction (up to 40% for RTX4090 and 20% for H100) of total GPU memory.
@@ 5.2 Experiment Setups
Table 3: Hardware specifications for our setups. In bandwidth, the number in the parentheses is the actual bandwidth we measured from our system.
(a) End-to-end latency speedup with Llama-3.2-3B with a single RTX4090 GPU.
(b) End-to-end latency speedup with Llama-3-8B with a single RTX4090 GPU.
(c) End-to-end latency speedup with Llama-3-8B with a single H100 GPU.
(d) End-to-end latency speedup with Llama-2-13B with a single H100 GPU.
@@
Figure 8: End-to-end latency speedup of TeleRAG and baseline on six RAG pipelines, three datasets, with H100 and RTX4090 GPU.
Metric Hardware nprobe
128 256 512
Precision - 95.9% 97.6% 98.9%
CPU Latency (ms) Desktop 89 176 349
Server 64 126 251
Table 4: Retrieval precision and baseline latency of FAISS for different nprobe values. Precision is measured by comparing top-10 results against greedy search without IVF.
5.3 Latency Evaluation
Baseline retrieval. Table 4 reports the retrieval precision and the baseline CPU retrieval latency of different nprobe. To conduct this test, we use 1024 random queries from NQ. For retrieval precision, we compare against the top-10 results from the greedy search. For the latency, we report the average retrieval time. Due to the limited scalability of the CPU, a higher nprobe directly leads to a longer retrieval latency even when using the state-of-the-art implementation from FAISS.
End-to-end latency. We evaluate the end-to-end latency of all the RAG pipelines and datasets under all the hardware settings described in §5.2. Our comprehensive end-to-end speedup results are presented in Figure 8.
We can observe that TeleRAG improves the execution speed consistently with various configurations. Across different pipelines, datasets, and model/hardware setups, the average speedup is 1.21x, 1.41x, and 1.72x for nprobe values of 128, 256, and 512, respectively. We observe a more significant speedup for larger nprobe, as increased nprobe settings incur longer retrieval latencies.
In the best case, we observe around 2.6x speedup (for Iter-RetGen, HotpotQA, Llama-3.2-3B, on RTX4090). This is because the Iter-RetGen pipeline employs a retrieval step in every iteration and has a short LLM output in general, providing a larger end-to-end speedup margin for accelerating the retrieval part.
Another pipeline that has a particularly high speedup is the SubQuestion pipeline, for which TeleRAG achieves around 2.5x speedup across all the datasets with Llama-3.2-3B on RTX4090. This pipeline uses LLM to generate multiple sub-questions and perform a batched retrieval with 3 to 4 generated queries at one time. The CPU retrieval performance of batched retrieval is significantly worse than that of GPU, as it has limited parallelism. In TeleRAG, we handle the batched retrieval efficiently by leveraging GPU’s high parallel computing capacity and, therefore, achieve a high speedup for the SubQuestion pipeline.
Overall, the performance gain is stable across different datasets, with less than 3% of variance across datasets on average. For a given GPU, TeleRAG has a higher speedup with a smaller LLM because retrieval time will account for a larger fraction of the end-to-end latency. Still, we have tangible speedup in RTX4090 with Llama-3-8B model (e.g., more than 1.5x when nprobe is 512, on average), even though prefetching memory budget is only 3.75GB (space left after holding LLM, embedding model, IVF cluster centroids and misc tensors on GPU) in this case.
Speedups on retrieval. Figure 9 shows the latency speedups on retrieval on NQ. We observe a consistent speedups for all the nprobe numbers but notice the greatest speedups of retrieval often occur at nprobe 256, with average speedups of 7.21x and 7.41x on RTX4090 and H100. This is because, with a larger nprobe, the retrieval performance of TeleRAG will start to be bounded by the missed clusters on the CPU (we have a fixed prefetch budget of different nprobe numbers). Nonetheless, TeleRAG still has a large latency improvement against the pure CPU baseline.
(a) On Desktop with a RTX4090 GPU.
(b) On Server with a H100 GPU.
Figure 9: Retrieval speedup on NQ.
(a) With Llama-3-8B.
(b) With Llama-2-13B.
@@ 5.3 Latency Evaluation
Figure 10: Latency breakdown for NQ on Server with a H100 GPU.
@@
HyDE 9 GB $77.23\%$ 7 GB $67.16\%$
SubQ 8 GB $63.26\%$ 7 GB $58.38\%$
Iter 5 GB $80.79\%$ 3 GB $61.46\%$
IRG 4 GB $44.49\%$ 2.5 GB $29.39\%$
FLARE 6 GB $58.17\%$ 3 GB $33.01\%$
S-RAG 3 GB $36.01\%$ 1.25 GB $14.98\%$
@@
Table 5: The prefetch budget and corresponding averaged cluster hit rate for each pipeline and hardware setup on NQ dataset. The target retrieval nprobe is 512.
5.4 Analysis
Latency breakdown and GPU compute interference. We further show the latency breakdown of running RAG pipelines with Llama-3-8B and Llama-2-13B on Server with H100, examining the case when the GPU’s memory capacity is enough for a good amount of prefetching. Figure 10 shows the breakdown of latency into LLM time and retrieval time. We compare the baseline’s retrieval latency (FAISS CPU) with TeleRAG’s retrieval latency in three different nprobe numbers.
The first thing we observe is that, although we set the prefetching budget to the amount that the data loading time can overlap with the pre-retrieval generation time, there is a slight slowdown in LLM generation time for TeleRAG. This is potentially because GPU’s compute performance will be impacted when there is a memory-loading workload, as noticed in  [103]. However, such impact is minimal, with less than 5% slowdown in general. We observe higher impacts for Iter-RetGen and Self-RAG with Llama-3-8B, which TeleRAG’s LLM geeration time is 28% and 20% longer than the baseline. It’s because these two pipelines have a relatively shorter LLM generation time, and thus, are more vulnerable to interference on GPU compute. However, with the significant amount of speedups on retrieval, TeleRAG achieves consistent speedups across all the pipelines despite having this small impact on the LLM generation.
@@ 5.4 Analysis
Prefetch budgets and cluster hit rates. Table 5 shows the prefetch budgets we set with the profiling-guided approach on RTX4090 and H100 for NQ. It also presents the averaged cluster hit rate achieved with this prefetch budget. From the table, we can see that it can generally attain a high cluster hit rate ($>50$%) for the cases where it can hold a larger amount of data for prefetching. For pipelines or scenarios that can hold for less than 3GB prefetching amount, we generally observe a relatively lower hit rate ($<35$%), limiting our benefits of reducing CPU’s search workloads. However, as observed from Figure 8, TeleRAG achieves from 1.2x to 1.5x end-to-end speedups for these pipelines. This is the combined benefit of reducing CPU workload and utilizing GPU to perform sorting on similarity distances, demonstrating the strong performance of TeleRAG across a wide range of scenarios.
Analysis on the cluster hit rate. We measure cluster hit rates across a wide range of prefetch amounts, examining scenarios with higher CPU-GPU bandwidth and sufficient GPU memory to accommodate larger prefetches. Our experiments are conducted on the HyDE pipeline with nprobe set to 512 and prefetching from 32 to 1024 clusters. We use nprobe as the prefetch metric (instead of bytes) to facilitate comparisons with optimal hit rates. We evaluate on NQ, HotpotQA, and TriviaQA, reporting average results in Figure 11.
We have three observations from Figure 11. (1) From $0\%$ to $50\%$ of nprobe, accuracy remains high, illustrating that the probability of hitting relevant clusters decreases by order of prefetching. (2) Between $50\%$ and $150\%$ of nprobe, coverage still increases rapidly but is slower than the first range. (3) After $150\%$ of nprobe, coverage plateaus at roughly $95\%$. These results suggest that prefetching the amount beyond the target nprobe is likely unnecessary, as the coverage increase is limited.
Figure 11: The cluster coverage rate for HyDE with different numbers of prefetched clusters. The target retrieval nprobe is $512$. Results are the average of NQ, HotpotQA and TriviaQA.
6 Related Work
6.1 Systems for RAG

### v2

@@ 4.4. Implementation Details
To enable concurrent CPU-GPU data transferring and LLM generation, we utilize the _copy(non_blocking=True) API in PyTorch and use separate CUDA streams for prefetching and LLM. In this way, the data copy operations will not block the GPU to perform computation, and thus the TeleRAG can do prefetching in parallel to the pre-retrieval LLM generation.
@@
#\# of chunks 21 million
#\# of IVF cluster 4096
Embed model Contriever (Izacard et al., 2021)
@@
Table 2. Detailed configurations of our retrieval index.
5. Evaluation
@@
We conduct extensive experiments to evaluate the effectiveness of TeleRAG. In this section, we describe the necessary details on how we set up the evaluations, present experimental results, and provide in-depth analysis and discussions.
5.1. Evaluation Datasets and RAG Models
@@
Datastore. We build a datastore based on the wiki_dpr dataset (Karpukhin et al., 2020), a popular dataset that contains 2.1 billion tokens from Wikipedia. Following previous works (Asai et al., 2023; Karpukhin et al., 2020; Min et al., 2023), we chunk the passages by every 100 tokens, and use Contriever (Izacard et al., 2021) to generate an embedding for each chunk. The embeddings have a hidden dimension of 768.
@@
Vector index. For the baseline retrieval, we build an IVF vector index using Faiss (Douze et al., 2024) on the datastore. As described in §4.4, we convert the Faiss index to a customized index in PyTorch for TeleRAG. See Table 2 for the detailed configurations of our vector index and datastore.
@@
LLMs. We evaluate TeleRAG on the Llama model family (Touvron et al., 2023) in three different sizes (Llama-3.2-3B, Llama-3-8B, Llama-2-13B) to represent different use cases.
@@
Table 3. Hardware specifications for our setups. In bandwidth, the number in the parentheses is the actual bandwidth we measured from our system.
RAG pipelines. We evaluate TeleRAG with six popular RAG pipelines. Figure 7 shows the overview of each pipeline. Note that although some pipelines lack pre-retrieval generation, the post-retrieval generation serves similar functionality for the retrieval of the next iteration. Below are brief descriptions of the evaluated RAG pipelines.
  1. (1)
    HyDE (Gao et al., 2023a) prompts LLM to generate a hypothetical paragraph and perform retrieval based on the embedding of the generated paragraph.
  2. (2)
    SubQestion (SubQ) [^3] Implemented in LlamaIndex as SubQuestionQueryEngine. (Liu, 2022) prompts LLM to generate multiple sub-questions and performs retrievals for each generated sub-question.
  3. (3)
    Iterative (Iter) [^4] Implemented in LlamaIndex as MultiStepQueryEngine. (Liu, 2022) prompts LLM to generate narrower questions first and iteratively refine them based on previous answers. At the end of each iteration, it prompts LLM to judge if the answer is good enough.
  4. (4)
    Iter-RetGen (IRG) (Shao et al., 2023) iteratively do retrieval and LLM generation for 3 iterations.
  5. (5)
    FLARE (Jiang et al., 2023c) iteratively issues retrievals based on the confidence (probability score) of predicted tokens for the upcoming sentence.
  6. (6)
    Self-RAG (S-RAG) [^5] We evaluate on the short-form version which only has one iteration. (Asai et al., 2023) uses the LLM to judge for retrieval, generate responses, and self-critique on the responses. We use fine-tuned models based on Llama2-7B and Llama2-13B from their official repository (Asai et al., 2024b).
Figure 7. Overview of six RAG pipelines that we evaluate.
Evaluation Datasets. We use three commonly used question-answering datasets, NQ (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), and TriviaQA (Joshi et al., 2017). For each dataset, we randomly sample 1024 queries and report the average unless otherwise specified.
5.2. Experiment Setups
Hardware setups. We evaluate TeleRAG on two hardware environments, Desktop and Server, which are equipped to represent the settings for the desktop and data center use cases. The Desktop has the RTX4090 (24 GB memory), and we use 3B and 8B models. The Server has the H100 (80 GB memory), and we use 8B and 13B models. These sizes represent common model sizes for RAG applications (Asai et al., 2023, 2024a). We do not evaluate the 13B model on Desktop as its model size (26 GB) exceeds RTX4090’s memory capacity. Table 3 summarizes the hardware configurations.
Nprobe and top-$k$. For the IVF index, nprobe is a hyperparameter that controls the trade-off between search efficiency and retrieval quality. A common heuristic is to set nprobe to $4\sqrt{N_{c}}$ (Zilliz, 2020). Given our index size of $N_{c}=4096$, we use nprobe = 256 ($4\sqrt{4096}$) by default, unless otherwise specified. For retrieval, top-$k$ denotes the number of most relevant documents to return. We use top-$k$ = 3, which is a standard choice in RAG.
RAG pipeline implementation. We implement the RAG pipelines with the FlashRAG framework (Jin et al., 2024b). For IRG, FLARE, and S-RAG, we use the framework’s default implementations. For the other pipelines, we reimplement them using FlashRAG’s APIs.
Benchmark methodology. We follow a benchmarking methodology from SGLang (Zheng et al., 2024) and use GPT-3.5-Turbo (OpenAI, [n. d.]) to run through each pipeline and record the input and output text of each step. During latency evaluation, we set the LLM’s output length based on the recorded text. This way, we ensure a fair latency comparison across different LLM models.
(a) End-to-end latency speedup with Llama-3.2-3B.
(b) End-to-end latency speedup with Llama-3-8B.
@@ 5.2. Experiment Setups
Figure 8. End-to-end latency speedup of TeleRAG and baseline on six RAG pipelines and three datasets, with RTX4090 GPU. nprobe is 256.
Baseline systems. To evaluate the latency of each pipeline, we construct a clean execution flow in Python that only contains LLM generation, datastore retrieval, and other necessary logical operations to fulfill each pipeline. For LLM generation, we use SGLang (Zheng et al., 2024), which is a state-of-the-art LLM inference engine. For retrieval, we use a popular retrieval library, Faiss (Douze et al., 2024), as the CPU-offloaded baseline. [^6] This baseline resembles frameworks like LlamaIndex and LangChain, which primarily function as wrappers around LLM serving and retrieval libraries. We build a system from the ground up to avoid the overhead introduced by their complex layers designed for advanced functionality.
Prefetching budget setups. Based on the methodology we described in §4.2, we profile each RAG pipeline with a small amount of 64 random samples from NQ (Kwiatkowski et al., 2019) and derive the prefetching budget of each pipeline.
(a) End-to-end throughput with six pipelines and different batch sizes using Llama-3-8B.
(b) End-to-end throughput with six pipelines and different batch sizes using Llama-2-13B.
Figure 9. End-to-end throughput of TeleRAG and the baseline across six RAG pipelines on the NQ dataset using Llama-3-8B and Llama-2-13B at different batch sizes on a H100 GPU. The nprobe is set to 256, and the $x$-axis represents the batch size.
Max prefetching memory. We set a maximum GPU memory limit for prefetching in each configuration. For Server, we allocate up to 16 GB. For Desktop, we allocate up to 10 GB and 3.75 GB for the 3B and 8B models, respectively. These settings demonstrate that TeleRAG can efficiently operate using only a small fraction (up to 40% for RTX4090 and 20% for H100) of total GPU memory.
5.3. Evaluation Results
Single-query latency on RTX4090. We evaluate the end-to-end RAG latency for single queries on Desktop equipped with RTX4090, representing a typical local scenario. Figure 8 shows the latency reduction of TeleRAG across three datasets and two LLMs (Llama-3.2-3B and Llama-3-8B).
As shown in Figure 8, TeleRAG consistently outperforms the baseline across all tested configurations. With Llama-3.2-3B, TeleRAG achieves average speedups of $1.55\times$, $1.54\times$, and $1.49\times$ on NQ, HotpotQA, and TriviaQA, respectively. The best speedup of $2.11\times$ is achieved in the Iter-RetGen pipeline on HotpotQA, because this pipeline involves frequent retrieval operations and generally short LLM outputs, which enhances the relative impact of retrieval acceleration.
Another notable improvement is in the SubQuestion pipeline, where TeleRAG achieves approximately $1.85\times$ speedup across all datasets. This pipeline uses LLM-generated sub-questions and performs batched retrievals of 3 to 4 queries simultaneously. CPU-based retrieval suffers from limited parallelism in such scenarios, but TeleRAG efficiently utilizes GPU parallelism, significantly enhancing performance.
When deploying Llama-3-8B, speedups with TeleRAG slightly decrease compared to Llama-3.2-3B, primarily due to increased LLM latency and reduced available memory for prefetching. Still, TeleRAG achieves approximately $1.3\times$ speedup across datasets, with a peak improvement of $1.82\times$ for Iter-RetGen pipeline on HotpotQA. Achieving these results with only $3.75$ GB of remaining GPU memory (after accounting for Llama-3-8B (16 GB), the embedding model (1 GB), and other miscellaneous tensors) highlights TeleRAG’s robust capability in accelerating RAG inference even under tight GPU memory constraints.
Figure 10. Simulated throughput [^8] Results for 1 GPU are real, not simulated.of TeleRAG on NQ dataset with different number of H100 GPUs.
Multi-query throughput on H100. To evaluate the TeleRAG’s performance in batched inference, we evaluate the end-to-end throughput on Server (H100) in batch sizes 1, 2, 4, and 8. We show the results of six evaluated RAG pipelines with Llama-3-8B and Llama-2-13B in Figure 9.
[…]
@@ 5.3. Evaluation Results
Figure 11. Latency breakdown for Llama-3-8B on NQ with a H100 GPU in different batch sizes. nprobe is 256.
Figure 12. Comparison of end‑to‑end retrieval latency for two batching strategies: naive mini‑batching and similarity‑aware greedy grouping. The Greedy Group bars are too short to be visible.
Multi-GPU Throughput. We also evaluate TeleRAG in a multi-GPU system to further show the scalability. For the evaluation, we construct a simulation framework around our single-GPU implementation. We first define a global batch size that the whole system will process at once, and then we construct mini-batches to distribute among each GPU. To form a mini-batch, we perform a greedy search based on the embedding similarity of queries, grouping similar queries into the same mini-batch to maximize the cluster hit rate. We present the simulated throughput of TeleRAG with a global batch size of 128 and mini-batch size of 4 with 1 to 8 GPUs. As shown in Figure 8, TeleRAG’s throughput scales well with the number of GPUs. When compared to 1 GPU results, the averaged throughput enhancements of TeleRAG are $1.7\times$, $3.1\times$, and $5.4\times$ with $2$, 44, and $8$ GPUs, respectively.
5.4. Analysis and Sensitivity Study
Latency breakdown. We further show the latency breakdown of running RAG pipelines with Llama-3-8B and Llama-2-13B on a single H100 GPU in different batch sizes in Figure 11. From Figure 11, we can observe that LLM latency grows sub-linearly with larger batch sizes. However, the latency for Faiss retrieval on CPU grows linearly with the batch size, dominating the overall latency when batch size is large. These results echo our findings in Figure 9, and show limited scalability of CPU retrieval in serving scenario. In contrast, TeleRAG significantly accelerates across all batch sizes and achieves a higher speedup from $1.4\times$ to $1.8\times$ when the batch size increases from 1 to 8.
Ablation on mini-batch strategies. In our multi-GPU results, we implement a greedy grouping strategy for mini-batching based on queries’ similarity in a global batch. Here, we examine the benefits and overhead of this strategy in Figure 12, where we show the end-to-end retrieval latency for mini-batch size $4$ with four GPUs and a global batch of 128 queries. The grouping overhead is minimal and the overall retrieval latency of greedy grouping mini-batching consistently outperform naive mini-batching across all pipelines.
(a) Llama-3.2-3B with a RTX4090 GPU.
(b) Llama-3-8B with a H100 GPU.
Figure 13. Single-query retrieval speedup on NQ with different nprobe values.
Retrieval speedups across nprobe. Figure 13 shows the retrieval latency reduction on NQ. We observe consistent speedups for all nprobe values. The greatest speedups are achieved at nprobe 256, with average speedups of $7.21\times$ and $7.41\times$ on RTX4090 and H100, respectively. As a larger nprobe value is used, the retrieval performance of TeleRAG becomes constrained by missed clusters on the CPU, given our fixed prefetch budget across varying nprobe values. However, RAG inference with higher nprobe will result in longer latency spent on the retrieval and hence, TeleRAG still has a significant latency improvement against the CPU retrieval baseline.
@@ 5.4. Analysis and Sensitivity Study
HyDE 9 GB $91.95\%$ 7 GB $87.42\%$
SubQ 8 GB $79.04\%$ 7 GB $76.38\%$
Iter 5 GB $95.51\%$ 3 GB $84.59\%$
IRG 4 GB $59.52\%$ 2.5 GB $50.34\%$
FLARE 6 GB $79.35\%$ 3 GB $56.67\%$
S-RAG 3 GB $71.29\%$ 1.25 GB $29.96\%$
@@
Table 4. The prefetch budget and corresponding averaged cluster hit rate for each pipeline and hardware setup on NQ dataset. The target retrieval nprobe is 256.
@@
Prefetch budgets and cluster hit rates. Table 4 shows the prefetch budgets we set with the profile-guided approach on RTX4090 and H100 for NQ. It also presents the average cluster hit rate achieved with this prefetch budget. From the table, we can see that it generally achieves a high cluster hit rate (¿5050%) when TeleRAG has a large prefetching budget. For cases where the budget is less than 2 GB, we observe a relatively lower hit rate (¡5050%), limiting the benefits of reducing the CPU’s search workloads. However, as observed from Figure 8, TeleRAG achieves from $1.2\times$ to $1.6\times$ end-to-end speedups for these pipelines, thanks to the combined benefit of reducing CPU workload and utilizing the GPU to perform sorting on similarity distances.
6. Related Work
6.1. Systems for RAG